There are 1,006 scouters joining the camping. If there are 206 more cub scouts than star scouts joining, how many star scouts are camping?: There are 1,006 scouters joining the camping. If there are 206 more cub scouts than star scouts joining, how many star scouts are camping?
Answer by jojo14344(326)

Let "x"= # of star scout
So, "x+206"=# of cub scouts
Then, x+x+206=1006
2x=1006-206
2x=800
x=400 -----------------> # of star scouts
x+206=400+206=606 -----> # of cub scouts
Total, # of star scouts + # of cub scouts =1006
400+606=1006
1006=1006
Thank you,
Jojo